OpenStudy (anonymous):

What is the slope of the line 5x + 2y = 10?

OpenStudy (anonymous):

rearrange this into slope-intercept form: y=mx+b. 2y = -5x + 10. y = -5/2x + 5. m = slope and b = y-intercept. The slope is -5/2. :)